OpenStudy (anonymous):

any body can derive the formula for a surface area and volume of cone through integration(calculus)??????????

OpenStudy (amistre64):

its been done already :)

OpenStudy (anonymous):

hahaha

OpenStudy (amistre64):

the typical radius is what.... is what, the slope of the linear between radius and height?

OpenStudy (amistre64):

you are adding up all the circumferences of the circles from 0 to h

OpenStudy (yuki):

volume is easy

OpenStudy (yuki):

the base of a cone is shaped as a circle

OpenStudy (amistre64):

[S] 2pi [f(x)] dx ...right?

OpenStudy (anonymous):

i tried integrating it bt i guess well have 2 use the volume of a frustum

OpenStudy (yuki):

so the base area is \[\pi r^2\]

OpenStudy (yuki):

if you use the disc method to integrate it into a cone, first we need to find the volume of each disc

OpenStudy (amistre64):

do we include the area of the base into surface area? or is it an open base?

OpenStudy (anonymous):

the disc method dint wrk out

OpenStudy (yuki):

if dx is the width of each disc, then the volume would be \[\pi r^2dx\]

OpenStudy (anonymous):

i tried this

OpenStudy (amistre64):

r = [f(x)] right?

OpenStudy (yuki):

now since the radius of a cone changes as the position of the disc changes, we need to find r in terms of x

OpenStudy (amistre64):

f(x) = h = linear equation from point (0,r) to (r,h)

OpenStudy (amistre64):

or... (0,x) to (x,y) :)

OpenStudy (amistre64):

ack...(x,0.. my mistake)

OpenStudy (anonymous):

dV= πr2^2dx

OpenStudy (amistre64):

him; deriving to get volume?

OpenStudy (amistre64):

integrate

OpenStudy (yuki):

if a cone has radius r and height h, the radius of the disc will be \[({-r \over h}x +r)\]

OpenStudy (anonymous):

r=xtant

OpenStudy (amistre64):

do we have a height and a radius?

OpenStudy (yuki):

so the volume of the disc is \[\pi ({-r \over h}x+r)^2dx\]

OpenStudy (yuki):

now we will find the limits of integration

OpenStudy (yuki):

this one is easy, the height is h, so the limit of integration is from x=0 to x=h thus the integral we are looking for to find the volume is \[\int\limits_{0}^{h} \pi ({-r \over h}x+r)^2dx\]

OpenStudy (yuki):

when you integrate it you will get \[{1 \over 3}\pi r^2\]

OpenStudy (anonymous):

dV=πr^2dh r=Rh/H dh=Hdr/R dV=πr^2Hdr/R V= πR^3 x H/3R

OpenStudy (yuki):

did it help at all ?

OpenStudy (anonymous):

yeah i think this takes care f d volume

OpenStudy (yuki):

now let's work on the surface area

OpenStudy (anonymous):

dS=2πrdl l=rL/R dl= Ldr/R S= 2Lπrdr/R πR^2 L/R = πRL

OpenStudy (anonymous):

DONE!!!!!!!!!!

OpenStudy (yuki):

the same idea is applied to the problem, but now instead of a volume, we are looking for an area. so instead of a disc, we will consider the lateral area of the disc, which can be found by \[2\pi r *dx\]

OpenStudy (anonymous):

thanks a lot fr the solution

OpenStudy (yuki):

again, the radius of a cone changes, so we need to find the radius in terms of x. But we already found it, which is again \[{-r \over h}x+r\]

OpenStudy (yuki):

again, the radius of a cone changes, so we need to find the radius in terms of x. But we already found it, which is again \[{-r \over h}x+r\]

OpenStudy (yuki):

so the lateral area of the discs can be found by \[2\pi[ {({-r \over h} )x+r}]dx\]

OpenStudy (yuki):

now for this one, you have to be careful with the limit of integration

OpenStudy (yuki):

the length of the slanted height is going to be my limit of integration because we are finding the lateral area of each discs so that all dx's add up to the slanted height.

OpenStudy (yuki):

so if we could find the slanted height we are good to go

OpenStudy (yuki):

now you can see that the slanted height makes a right triangle with the height of the cone and the radius of the base

OpenStudy (yuki):

so if I call L my slanted height, \[L = \sqrt{h^2 + r^2}\]

OpenStudy (anonymous):

S.A = L*2pi*r, where L is the slant length V = A*h, in this case height is dx, r = x so we have; \[\pi\int_0^x x^2 dx = (1/3)\pi x^3 \]

OpenStudy (yuki):

so the integration we have to do is \[\int\limits_{0}^{L} 2\pi[({-r \over h})x+r]dx\]

OpenStudy (yuki):

which ends up being \[\pi rL\]

OpenStudy (yuki):

or \[\pi r \sqrt{h^2+r^2}\]

OpenStudy (yuki):

I hope that helps :)

OpenStudy (yuki):

Give me about 5min to get the picture of why the radius is \[{-r \over h}x+r\]

OpenStudy (anonymous):

dV=πr^2dh r=Rh/H dh=Hdr/R dV=πr^2Hdr/R V= πR^3 x H/3R

OpenStudy (anonymous):

thats the vol

OpenStudy (yuki):

It would be nice if you could tell me if it helped or not ... :)

OpenStudy (anonymous):

ya it helped thanks